A 6 FT. TALL MAN IS STANDING NEAR A TREE ON LEVEL GROUND. HIS SHADOW IS 4FT LONG. IF THE LENGTH OF THE TREE'S SHADOW IS EQUAL TO 18 FT.- HOW TALL IS THE TREE? ** THE ANSWER IS 27FT- I
AM CONFUSED ON HOW THEY GOT THAT ANSWER. PLEASE HELP. THANK YOU, CHRISTINA

Answer by stanbon(29565)

A 6 FT. TALL MAN IS STANDING NEAR A TREE ON LEVEL GROUND. HIS SHADOW IS 4FT LONG. IF THE LENGTH OF THE TREE'S SHADOW IS EQUAL TO 18 FT.- HOW TALL IS THE TREE?
-----------
It's a proportion.
object/shadow = object/shadow
Your Problem:
x/18 = 6/4
x = 18(3/2)
x= 27 ft.
============
Cheers,
Stan H>